the number of neutrons present in 44 g 14CD4 is

Grade:12th pass

2 Answers

Bharat Khanna
10 years ago

Molar mass of CD4=14+2*4=22

No. of moles of CD4=44/22=2

No. of neutron in 1 mole CD4=(8+4)*Na

=12*Na(Na is avogadro''s number, i.e. 6.022*10^23)

No. of neutron in 2 mole =24*Na

Vikash Kumar
10 years ago

no. of moles= 44/(14+8) = 2

1 14C has 8 neutrons

1 2D has 2 neutrons

1 CD4 molecule has 16 neutrons.

2 moles will have = 16 * 6.022 * 1023  neutrons. = 9.635 * 1024 

Thanks & Regards
Vikash Kumar, 
askIITians Faculty
B.Tech, IIT Delhi